The sum of two consecutive even integers is -298. what are the integers?
Answer by stanbon(75887)

The sum of two consecutive even integers is -298. what are the integers?
---
1st: 2x
2nd: 2x+2
-----
Equation:
sum = -298
2x + 2x+2 = -298
4x+2 = -298
4x = -300
x = -75
-----
1st: 2x = -150
2nd: 2x+2 = -148
======================
Cheers,
Stan H.